john wants to make a triangular garden in his back yard. which of the following are possible dimensions?
A. 4FT*5ft*10ft
B. 6ft*6ft*12ft
C. 6ft*8ft*10ft
D. 8ft*12ft*20ft

Answer by JBarnum(2146)

the largest side cant be greater than or equal to the combined total of the 2 lesser sides
A. 4,5,10
4+5=9
10 is greater then 9 so not a triangle
B. 6,6,12
6+6=12
12 = 12 and cant be a triagle as it would be the exact same line
D. 8,12,20
8+12=20
20=20
C. 6,8,10
6+8=14
10 is less than 14 so this is a triangle